# Discrete Spatial Diffusion

### v1

@@ Abstract
Generative diffusion models have achieved remarkable success in producing high-quality images. However, because these models typically operate in continuous intensity spaces—diffusing independently per pixel and color channel—they are fundamentally ill-suited for applications where quantities such as particle counts or material units are inherently discrete and governed by strict conservation laws like mass preservation, which limits their applicability in scientific workflows. To address this limitation, we propose Discrete Spatial Diffusion (DSD), a framework based on a continuous-time, discrete-state jump stochastic process that operates directly in discrete spatial domains while strictly preserving mass in both forward and reverse diffusion processes. By using spatial diffusion to achieve mass preservation, we introduce stochasticity naturally through a discrete formulation. We demonstrate the expressive flexibility of DSD by performing image synthesis, class conditioning, and image inpainting across widely-used image benchmarks, with the ability to condition on image intensity. Additionally, we highlight its applicability to domain-specific scientific data for materials microstructure, bridging the gap between diffusion models and mass-conditioned scientific applications.
Discrete diffusion, hard constraints, intensity preserving, diffusion models, generative models, digital rocks, batteries, machine learning, geometry generation, porosity, volume fraction, preserving, genAI, material generation
@@ 1 Introduction
Diffusion-based generative models have emerged as powerful tools for high-quality image generation (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b). Typically, these models inject noise into the images, then learn to reverse this noise-adding process to recover meaningful structure. In most frameworks, this is based on an Itô Stochastic Differential Equation (SDE) with Gaussian noise. While effective for many vision tasks, these approaches inherently assume continuous pixel intensities, which can cause difficulty when dealing with the discrete nature of many datasets. Beyond vision-related tasks, such as in the physical sciences, there are many applications which require discrete physical quantities, such as particle counts in a simulation, or phases in materials microstructure. Conservation of total quantities (such as mass) can be critical for scientific applications, and so generative modeling which can operate under constrained, discrete pixel intensities would be scientifically valuable. Such a capability might also prove useful within vision tasks, such as inpainting and super-resolution.
@@
Scientific and engineering studies of the natural world using computational techniques often involve discrete variables in space and/or time. On microscopic scales, everyday materials exhibit extremely complex structural patterns which encode the history of their formation, and play a large role in how the material functions on a macroscopic level. An important and wide-reaching field of study is materials microstructure, which is used in materials design (Gu et al., 2018), forensic analysis, hydrology (Blunt et al., 2013), energy storage (Simon & Gogotsi, 2008), and even medicine, such as in bone structure studies (Montoya et al., 2021). For example, crystal grain shapes can give rise to complex stress patterns which affect the yield strength of a metal (Calcagnotto et al., 2011). A materials microstructure can often be represented in terms of a small number of discrete phases which describe the underlying chemical structures involved in the microstructure. In sandstone, the overall arrangement of nanocrystals is highly disordered and gives rise to complex pore structures, through which subsurface water flows, and this microstructure can have an enormous influence on the rate of transport of fluids and contaminants. Microstructure of electrodes is also known to have an immense impact on the characteristics of electrochemical devices (Phogat et al., 2024). Small changes in thermodynamic properties can cause drastic changes in microstructure, such as in stainless steels (Xiong et al., 2010), necessitating study of microstructure as a function of phase contents. Furthermore, gathering real-world data on these systems is often complex and expensive; decades of work have been applied to computational modeling of the generation and consequences of microstructure prior to the widespread popularization of machine learning (Torquato, 2002).
@@
In this work, we introduce Discrete Spatial Diffusion (DSD), a discrete-state Markov chain-based diffusion framework in which the forward process redistributes discrete units of intensity in space. Unlike previous diffusion models, DSD exactly preserves total intensity throughout both the forward and reverse phases, ensuring that global properties—such as mass fractions or total particle count—are exactly conserved. We demonstrate that DSD not only enables scientific applications but also applies to more conventional tasks, like image generation and in-painting in discrete domains. By directly modeling discrete transitions, DSD paves the way for generative modeling under mass conservation, allowing models that specialize for constrained conditions in scientific applications and beyond.
@@ 2.1 Related work
Among the body of literature on generative diffusion models, originating from the pioneering work of Sohl-Dickstein et al. (2015), the most relevant to our work fall into two broad categories: (1) those employing discrete-state Markov chains to introduce noise in the forward process (Hoogeboom et al., 2021; Austin et al., 2021; Campbell et al., 2022; Santos et al., 2023; Sun et al., 2022; Lou et al., 2024), and (2) those incorporating spatial dynamics into the forward diffusion process (Bansal et al., 2022; Rissanen et al., 2022; Hoogeboom et al., 2021).
@@
Generative diffusion modeling based on discrete-state Markov chains has become an active area of research in recent years. Early work, such as Hoogeboom et al. (2021); Austin et al. (2021), introduced discrete-state and discrete-time Markov chains as an alternative to the Gaussian noise used in conventional diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b). Campbell et al. (2022) generalized these formulations to a continuous-time framework, providing a more rigorous theoretical foundation for discrete-state generative diffusion modeling. Santos et al. (2023) employed operator algebraic analysis to formally establish the existence of the reverse-time dynamics and derived the stochastic generator for arbitrary discrete-state Markov processes. Similar formulations were independently developed by Sun et al. (2022) and Lou et al. (2024), with an emphasis on defining and estimating score functions for discrete-state systems. The Markov process operates in intensity space in all the aforementioned diffusion models, treating each pixel as an independent stochastic process (Fig. 1(a): Gaussian; Fig. 1(b): Discrete).
@@
This study focuses on a spatially correlated process for generative modeling for two reasons: (1) for structured images, a more natural approach is to incorporate spatial correlations into the generative process, and (2) spatially decorrelated noise makes it difficult to preserve total intensity. A spatially correlated approach has been explored for continuous systems. Cold Diffusion (Bansal et al., 2022) introduced a deterministic blurring transformation, where image degradation follows a predefined forward process, and reconstruction is learned as an inverse mapping. However, lacking a probabilistic latent distribution (as in VAEs (Kingma & Welling, 2014)), Cold Diffusion is not a true generative model. Inverse Heat Dissipation Model (IHDM, Rissanen et al. (2022)) uses the heat equation as a corruption model. Since the heat equation is deterministic and reversible (except for the homogeneous solution at $t\rightarrow\infty$ is singular), a naïve inversion would again result in deterministic reconstructions. Uncorrelated Gaussian noise was added to the heat equation to overcome this limitation, relaxing the deterministic process into a probabilistic Itô diffusion. Later, Blurring Diffusion Model (BDM, Hoogeboom & Salimans (2022)) recognized that IHDM could be recast as a Gaussian diffusion model in the spectral domain. BDA extended IHDM and achieved SOTA generative performance, validating the hypothesis that spatially structured diffusion processes can enhance image generation. Nevertheless, the probabilistic formulation of IHD and BDM only preserves mass on average, not exactly per-sample, and their continuous-state nature makes it difficult to apply to discrete datasets.
@@
Figure 1: Schematic diagrams of various diffusion models. (a) Gaussian Diffusion relies on the Ornstein–Uhlenbeck diffusion process in the intensity space.(b) Previous discrete-state diffusion models rely a discrete-state Markov process of the transition of discretized intensities. (c) Discrete Spatial Diffusion (this study) relies on a Markov jump process of the intensity units on the discrete lattice, conserving the total intensity seperately in each color channel.
@@
Our goal of generating samples with exactly conditioned total intensity aligns with conditional diffusion modeling. However, existing approaches all rely on some degree of approximation. Song & Ermon (2019) proposed a simple conditional sampling method by passing class labels into the neural network during training, but this does not guarantee exact enforcement of the condition in generated samples. A more structured approach was introduced by Chung & Ye (2022); Chung et al. (2022c, b), which interleaved projection steps with diffusion sampling to enforce linear constraints in image generation. However, these projections disrupt the exactness of the forward corruption and reverse inference dynamics (Anderson, 1982; Campbell et al., 2022; Santos et al., 2023), leading to a mismatch between the projected and true data manifolds. To address this, Chung et al. (2022a) eliminated projection steps but instead relaxed deterministic constraints into a probabilistic formulation via a noisy measurement model. However, this method does not apply to deterministic constraints, as it becomes singular in the limit of zero measurement noise. An alternative approach leverages Bayes’ theorem for a posteriori conditional sampling, that is, $p(\text{S}|\text{C})\propto p(\text{S})p(\text{C}|\text{S})$, where “S” stands for samples and “C” for condition(s). Because $p(\text{S})$ is given by a trained unconditional diffusion model, conditioning can be performed if one has $p(\text{C}|\text{S})$, which is however intractable for arbitrary data distributions [^1] It is challenging because the constraint is imposed on the final samples at the end of the inference, but the conditioning “S” are samples generated during the inference.. Existing methods approximate this term crudely or by training a separate classifier as in Song et al. (2021b), or by a Gaussian approximation with moment-matching as in (Finzi et al., 2023; Du et al., 2024). None of these methods guarantees the generated samples are exactly conditioned.
2.2 Summary of our contributions
@@
The principal contribution of our work is that it provides a new capability for diffusion models to preserve intensity exactly in a fully discrete-state context. The approach is based entirely on how the diffusion process is built, and how the model is trained; it is readily usable with existing diffusion model neural network (NN) architectures. To our knowledge, this is the first diffusion model to incorporate a spatially correlated noise, which is accomplished using a stochastic jump process allow units of intensity to perform a random walk. This fact also demonstrates that more complex noise processes can themselves be tractable. We furthermore demonstrate that such a model is powerful enough for conventional image synthesis tasks. The relevance and power of the approach is then demonstrated through application to scientific data in the field of materials microstructure, where the ability to generate complex data-driven images with constrained total intensity is highly desirable.
@@ 3.1 Corruption Process
In this manuscript, we adopt the language of image processing and consider 2-dimensional images, although the context and spatial dimensionality of the data are not constrained by the mathematical framework provided here. We treat a digital image with discretized intensities $I_{x,y,c}\in\mathbb{Z}_{\geq 0}$ at pixel $(x,y)\in\{1,\ldots,W\}\times\{1,\ldots,H\}$ in color channel $c\in\{1,\ldots,C\}$. Within the DSD framework, the image is treated as a spatially organized collection of particles; one for each intensity unit. Below, we will interchangeably use “particles” and “intensity units” to denote these fundamentally discrete units. Specifically, $I_{x,y,c}=n$ implies $n$ particles of type $c$ at location (x,y), and the total number of particles of the system is $\sum_{x=1}^{H}\sum_{y=1}^{W}\sum_{c=1}^{C}I_{x,y,c}$. In the forward stochastic process with the time parameter $t$, each of the particles in the system independently performs a continuous-time and discrete-state random walk:
@@
where $r$ is the transition rate of the particle jumping to one of their nearest neighbors, and $\nu$ is a set of four directions the particles can hop to their nearest neighbors. Note that the particles perform jumps in the $(x,y)$ space at random times, but do not change their color coordinate $c$. A schematic diagram is shown in Fig. 1c. We impose either no-flux boundary condition, such that the transition rates of jumping out of the image domain are zeros, or periodic boundary conditions so that a jump to $x=W+1$ becomes a jump to $x=0$, vice-versa, and analogously for $y$. Note that the forward process conserves the total number of particles, $\sum_{x=1}^{H}\sum_{y=1}^{W}I_{x,y,z}$ in each color channel independently.
We refer to the spatial hopping process (1) as the Discrete Spatial Diffusion (DSD), noting the “discreteness” refers to both the discretized intensity units and the discreteness of the spatial lattice $\{1,\ldots,W\}\times\{1,\ldots,H\}$ where the particles are allowed to reside. DSD, as well as similar discrete-state random walks, have been extensively studied in non-equilibrium statistical physics and stochastic processes (Van Kampen (2007), Gardiner (2009), Giuggioli (2020) and references therein). The evolution of the probability distribution of the single random walk in the continuum space limit, under the appropriate scaling of the transition rate (Einstein, 1905), converges to the Fokker–Planck Equation (FPE, Van Kampen (2007); Risken (1984)), which is mathematically identical to the heat equation. Because of the duality between the probabilistic FPE and the deterministic heat equation (Lawler, 2010), DSD can be considered as a microscopic description of the macroscopic heat dissipation that inspires IHD and BDM. Notably, the correlated noise is built in DSD, in contrast to the heuristic addition of uncorrelated Gaussian noise in IHD and BDM. Figure 2 illustrates the application of DSD to a sample image. Due to the stochasticity of the random jumps, the limiting behavior ($t\rightarrow\infty$) of this process is a random configuration with no discernible structure or similarity to the original spatial organization aside from the conserved global particle counts in each color channel.
@@
The forward solution, the transition probabilities $p_{t}(x,y,c|x_{0},y_{0},c_{0}):=\mathbb{P}\{X_{t}=x,Y_{t}=y,Z_{t}=z|X_{0}=x_{0},Y_{0}=c_{0},C_{0}=c_{0}\}$, can be computed by integrating the Master Equation (Van Kampen, 2007; Gardiner, 2009; Weber & Frey, 2017). This corresponds to exponentiating the Markov transition matrix of the process defined in Eq. (1). While the matrix exponential required numerically for no-flux boundaries is expensive, the solution can be stored and reused to generate corrupted images and to compute the reverse-transition rates (see Sec. 3.3) for learning. When periodic boundary conditions are imposed, the transition matrix is diagonal in the discrete Fourier space, facilitating the efficient computation of $p_{t}(\cdot|\cdot)$ (see Appendix C).

### v2

@@ Abstract
Generative diffusion models have achieved remarkable success in producing high-quality images. However, these models typically operate in continuous intensity spaces, diffusing independently across pixels and color channels. As a result, they are fundamentally ill-suited for applications involving inherently discrete quantities—such as particle counts or material units—that are constrained by strict conservation laws like mass conservation, limiting their applicability in scientific workflows. To address this limitation, we propose Discrete Spatial Diffusion (DSD), a framework based on a continuous-time, discrete-state jump stochastic process that operates directly in discrete spatial domains while strictly preserving particle counts in both forward and reverse diffusion processes. By using spatial diffusion to achieve particle conservation, we introduce stochasticity naturally through a discrete formulation. We demonstrate the expressive flexibility of DSD by performing image synthesis, class conditioning, and image inpainting across standard image benchmarks, while exactly conditioning total image intensity. We validate DSD on two challenging scientific applications: porous rock microstructures and lithium-ion battery electrodes, demonstrating its ability to generate structurally realistic samples under strict mass conservation constraints, with quantitative evaluation using state-of-the-art metrics for transport and electrochemical performance.
@@ 1 Introduction
Diffusion-based generative models have emerged as powerful tools for high-quality image generation [66, 36, 69]. Typically, these models inject noise into the images, then learn to reverse this noise-adding process to recover meaningful structure. In most frameworks, this is based on an Itô Stochastic Differential Equation (SDE) with Gaussian noise. While effective for many vision tasks, these approaches inherently assume continuous pixel intensities, which can cause difficulty when dealing with the discrete nature of many datasets. Tasks beyond vision, such as those in the physical sciences have many applications which require discrete physical quantities, such as particle counts in a simulation, or phases in materials microstructure. Conservation of total quantities is critical for scientific applications, and so generative modeling which can operate under constrained, discrete pixel intensities would enable scientifically grounded and physically consistent synthesis. Such a capability might also prove useful within vision tasks, such as inpainting and super-resolution.
@@
Scientific and engineering studies of the natural world using computational techniques often involve discrete variables in space and/or time. On microscopic scales, everyday materials exhibit extremely complex structural patterns which encode the history of their formation, and play a large role in how the material functions on a macroscopic level. An important and wide-reaching field of study is materials microstructure, which is used in materials design [35], forensic analysis, hydrology [7], energy storage [65], and even medicine, such as in studies of bone structure [53]. For example, crystal grain shapes can give rise to complex stress patterns which affect the yield strength of a metal [10]. Materials microstructures are frequently represented by a small number of discrete phases that describe their underlying chemical structures. In sandstone, for example, the overall arrangement of nanocrystals is highly disordered, and gives rise to complex pore structures, through which subsurface water flows. This microstructure has an enormous influence on the rate of transport of fluids and contaminants [6]. Microstructure of electrodes is also known to have an immense impact on the characteristics of electrochemical devices [58]. Small changes in thermodynamic properties can cause drastic changes in microstructure, such as in stainless steels [79], requiring the study of microstructure as a function of phase contents. Furthermore, gathering real-world data is often complex and expensive; decades of work have been applied to computational modeling of the generation and consequences of microstructure prior to the widespread popularization of machine learning [72]. While diffusion models have been constructed for microstructure [52, 37, 47, 46, 80, 4, 25], they have not handled exact porosity (intensity) constraints.
@@
In this work, we introduce Discrete Spatial Diffusion (DSD), a discrete-state Markov chain-based diffusion framework in which the forward process redistributes discrete units of intensity in space (Fig. 1c). Unlike previous diffusion models, DSD exactly preserves total intensity throughout both the forward and reverse phases, ensuring that global properties—such as mass fractions—are exactly conserved by representing them in terms of conserved particles. We demonstrate that DSD enables scientific applications and can extend the capabilities of conventional image processing tasks, like image generation and inpainting in discrete domains. By directly modeling discrete transitions, DSD enables generative modeling under conservation laws, allowing models that specialize for constrained conditions in scientific applications and beyond.
@@ 2 Background
Among the body of literature on generative diffusion models, originating from the pioneering work of Sohl-Dickstein et al. [66], the most relevant to our work fall into two broad categories: (1) those employing discrete-state Markov chains to introduce noise in the forward process [39, 3, 11, 63, 70, 51, 64], and (2) those incorporating spatial dynamics into the forward diffusion process [5, 61, 39].
@@
Generative diffusion modeling based on discrete-state Markov chains has become an active area of research in recent years. Early work, such as Hoogeboom et al. [39], Austin et al. [3], introduced discrete-state and discrete-time Markov chains as an alternative to the Gaussian noise used in conventional diffusion models [66, 36, 69]. Campbell et al. [11] generalized these formulations to a continuous-time framework, providing a more rigorous theoretical foundation for discrete-state generative diffusion modeling. Santos et al. [63] employed operator algebraic analysis to formally establish the existence of the reverse-time dynamics and derived the stochastic generator for arbitrary discrete-state Markov processes. Similar formulations were independently developed by Sun et al. [70] and Lou et al. [51], with an emphasis on defining and estimating score functions for discrete-state systems. The Markov process operates in intensity space in all the aforementioned diffusion models, treating each pixel as an independent stochastic process (Fig. 1(a): Gaussian; Fig. 1(b): Discrete).
@@
This study focuses on a spatially correlated process for generative modeling for two reasons: (1) for structured images, it is more natural to incorporate spatial correlations into the generative process, and (2) spatially decorrelated noise makes it difficult to preserve total intensity. A spatially correlated approach has been explored for continuous systems. Cold Diffusion [5] introduced a deterministic blurring transformation, where image degradation follows a predefined forward process, and reconstruction is learned as an inverse mapping. However, lacking a probabilistic latent distribution (as in VAEs [42]), Cold Diffusion is not a true generative model. Inverse Heat Dissipation Model (IHDM, Rissanen et al. [61]) uses the heat equation as a corruption model. Since the heat equation is deterministic and reversible (except that the homogeneous solution at $t\rightarrow\infty$ is singular), a naïve inversion would again result in deterministic reconstructions. Uncorrelated Gaussian noise was added to the heat equation to overcome this limitation, relaxing the deterministic process into a probabilistic Itô diffusion. Later, Blurring Diffusion Model (BDM, Hoogeboom and Salimans [38]) recognized that IHDM could be recast as a Gaussian diffusion model in the spectral domain. BDA extended IHDM and achieved SOTA generative performance, validating the hypothesis that spatially structured diffusion processes can enhance image generation. Nevertheless, the probabilistic formulation of IHD and BDM only preserves intensity on average, not exactly per-sample, and their continuous-state nature makes it difficult to apply to discrete datasets.
@@
Figure 1: Schematic diagrams illustrating how intensity is modeled in different diffusion frameworks. (a) Gaussian Diffusion relies on the Ornstein–Uhlenbeck process in continuous intensity space. (b) Prior discrete-state diffusion models apply a discrete-state Markov process to independent pixel intensities. (c) Discrete Spatial Diffusion (this work) relies on a Markov jump process of the intensity units over a discrete spatial lattice, redistributing particles while exactly conserving total intensity per color channel.
@@
Our goal of generating samples with exactly conditioned total intensity aligns with conditional diffusion modeling. However, existing approaches all rely on some degree of approximation. Song and Ermon [68] proposed a simple conditional sampling method by passing class labels into the neural network during training, but this does not guarantee exact enforcement of the condition in generated samples. A more structured approach was introduced by Chung and Ye [15], Chung et al. [18, 17], which interleaved projection steps with diffusion sampling to enforce linear constraints in image generation. However, these projections disrupt the exactness of the forward corruption and reverse inference dynamics [2, 11, 63], leading to a mismatch between the projected and true data manifolds. To address this, Chung et al. [16] eliminated projection steps but instead relaxed deterministic constraints into a probabilistic formulation via a noisy measurement model. However, this method does not apply to deterministic constraints, as it becomes singular in the limit of zero measurement noise. An alternative approach leverages Bayes’ theorem for a posteriori conditional sampling, that is, $p(\text{S}|\text{C})\propto p(\text{S})p(\text{C}|\text{S})$, where “S” stands for samples and “C” for condition(s). Because $p(\text{S})$ is given by a trained unconditional diffusion model, conditioning can be performed if one has $p(\text{C}|\text{S})$, which is however intractable for arbitrary data distributions [^1] It is challenging because the constraint is imposed on the final samples at the end of the inference, but the conditioning “S” are samples generated during the inference.. Existing methods approximate this term crudely or by training a separate classifier as in Song et al. [69], or by a Gaussian approximation with moment-matching as in [27, 23]. None of these methods guarantees that the generated samples are exactly conditioned.
@@
The principal contribution of our work is that it provides a new capability for diffusion models to preserve intensity exactly in a fully discrete-state context. The approach is based entirely on how the diffusion process is built, and how the model is trained; it is readily usable with existing diffusion model neural network (NN) architectures. To our knowledge, this is the first diffusion model to incorporate spatially correlated noise, achieved through a stochastic jump process that allows units of intensity to perform a random walk. This fact also demonstrates that more complex noise processes can themselves be tractable. We furthermore demonstrate that such a model is effective for conventional image synthesis tasks. The relevance and power of the approach is then demonstrated through application to scientific data in the field of materials microstructure, where the ability to generate complex data-driven images with constrained total intensity is key to physically meaningful generative modeling and scientific reliability.
@@ 3.1 Corruption Process
In this manuscript, we adopt the language of image processing and consider 2-dimensional images, although the application context and spatial dimensionality of the data are not constrained by the mathematical framework provided here. We treat a digital image with discretized intensity values $I_{x,y,c}\in\mathbb{Z}_{\geq 0}$ at pixel $(x,y)\in\{1,\ldots,W\}\times\{1,\ldots,H\}$ in color channel $c\in\{1,\ldots,C\}$. Within the DSD framework, the intensity values in the image are treated as a spatially organized collection of particles, with one particle for each intensity unit. Below, we will interchangeably use “particles” and “intensity units” to denote these fundamentally discrete units. Specifically, $I_{x,y,c}=n$ implies $n$ particles of type $c$ at location (x,y), and the total number of particles of the system is $\sum_{x=1}^{H}\sum_{y=1}^{W}\sum_{c=1}^{C}I_{x,y,c}$. In the forward stochastic process with the time parameter $t$, each of the particles in the system independently performs a continuous-time and discrete-state random walk:
@@
where $r$ is the transition rate of the particle jumping to one of their nearest neighbors, and $\nu$ is a set of four directions the particles can hop to their nearest neighbors. A schematic diagram is shown in Fig. 1c. Note that the particles perform jumps in the $(x,y)$ space at random times, but do not change their color coordinate $c$. Because of this, the forward process conserves the total number of particles $\sum_{x=1}^{H}\sum_{y=1}^{W}I_{x,y,z}$ in each color channel independently. We impose either no-flux boundary condition, such that the transition rate for jumps out of the image domain are zero, or periodic boundary conditions, so that a jump to $x=W+1$ becomes a jump to $x=0$, vice-versa, and analogously for $y$.
We refer to the spatial hopping process (Eq. (1)) as the Discrete Spatial Diffusion (DSD), noting the “discreteness” refers to both the discretized intensity units and the discreteness of the spatial lattice $\{1,\ldots,W\}\times\{1,\ldots,H\}$ where the particles are allowed to reside. DSD, as well as similar discrete-state random walks, have been extensively studied in non-equilibrium statistical physics and stochastic processes (Van Kampen [75], Gardiner [28], Giuggioli [33] and references therein). The evolution of the probability distribution of the single random walk in the continuum space limit, under the appropriate scaling of the transition rate [26], converges to the Fokker–Planck Equation (FPE, Van Kampen [75], Risken [60]), which is mathematically identical to the heat equation. Because of the duality between the probabilistic FPE and the deterministic heat equation [44], DSD can be considered as a microscopic description of the macroscopic heat dissipation that inspires IHD and BDM. Notably, the correlated noise is built in DSD, in contrast to the heuristic addition of uncorrelated Gaussian noise in IHD and BDM. Fig. 2 illustrates the application of DSD to a sample image. Due to the stochasticity of the random jumps, the limiting behavior ($t\rightarrow\infty$) of this process is a random configuration with no discernible structure or similarity to the original spatial organization aside from the conserved global particle counts in each color channel.
@@
The forward solution and the transition probabilities $p_{t}(x,y,c|x_{0},y_{0},c_{0}):=\mathbb{P}\{X_{t}=x,Y_{t}=y,Z_{t}=z|X_{0}=x_{0},Y_{0}=c_{0},C_{0}=c_{0}\}$, can be computed by integrating the Master Equation [75, 28, 77]. This corresponds to exponentiating the Markov transition matrix of the process defined in Eq. (1). While the matrix exponential required numerically for no-flux boundaries is expensive, the solution can be stored and reused to corrupt images and to compute the reverse-transition rates (see Sec. 3.3) for learning. When periodic boundary conditions are imposed, the transition matrix is diagonal in the discrete Fourier space, facilitating the efficient computation of $p_{t}(\cdot|\cdot)$ (see Sec. C).